### v1

@@ Abstract
Generative machine learning models have begun to transform protein engineering, yet no principled framework for conditioning on auxiliary information in a plug-and-play manner exists; one may want to iteratively incorporate experimental feedback, or make use of an existing classifier—such as for predicting enzyme commission number—in order to guide the sampling of the generative model to generate sequences with desired properties. Herein, we present ProteinGuide, a rigorous and general framework to achieve just that: through unifying a broad class of protein generative models that includes masked language, (order-agnostic) autoregressive, diffusion and flow-matching models, we provide an approach to statistically condition pre-trained protein generative models. We demonstrate applicability of our approach by guiding each of two commonly used protein generative models, ProteinMPNN and ESM3, to generate amino acid and structure token sequences conditioned on several user-specified properties, namely, enhanced stability and CATH-labeled fold generation.
@@
Generative models are beginning to better enable protein engineers to create bespoke proteins for therapeutic, agricultural and environmental uses [1, 2, 3]. The upshot of any AI-based protein engineering approach is to ultimately specify a protein sequence, or a set of such sequences, even when designing by way of structure. For example, one might first use a backbone structure generative model such as RFdiffusion [2], Chroma [3], or ESM3 [4], followed by use of a backbone-conditioned sequence generative model (also called an inverse-folding model), such as ProteinMPNN [1] or ESM-IF1 [5], to generate sequences likely to fold into the specified backbones. Alternatively, one might forgo generating a backbone structure intermediary and instead generate sequences directly from an “evolutionary” sequence-only model such as ESM3 [4], ProGen [6], EvoDiff [7], DPLM [8] or many others [9, 10]. In all cases, the final step employs a generative model of sequences.
@@
A key ingredient missing from the current protein engineering AI menu is the plug-and-play ability to post hoc “guide” these sequence generative models toward properties of interest, because the technical underpinnings to do so have remained elusive. By guide, we mean, formally, to condition the probability distribution encoded by an existing generative model, $p(x)$, on newly-desired properties of interest, $p(y\in Y|x)$, so as to be able to generate (i.e., sample) desired sequences $x\sim p(x|y\in Y)$. Importantly, we want to do this conditioning in a plug-and-play manner, namely, without re-training the existing model. Although it is currently possible to re-train generative models to achieve property conditioning [11, 12], the plug-and-play characteristics of guidance would enable everyday users to readily adapt existing large, expensive-to-train models of interest to suit their needs, on-the-fly [13, 14, 15].
@@
Figure 1: Equivalence between classes of generative models enables guidance with ProteinGuide. a. Schematic showing how masked language models (e.g., ESM3) and order-agnostic autoregressive models (e.g., ProteinMPNN) can be viewed through the lens of masking diffusion/flow models. This re-interpretation enables the application of predictor-guidance to obtain ProteinGuide. Masking of one sequence position is denoted with the symbol ”?”, and colored dots help to visually tag individual sequences. The sequence being masked during training is M K R S V T; $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model. The shared training objective involves randomly masking out parts of the sequence and then predicting the masked sequence given the remaining unmasked context b. Contrasting unguided generation with predictor-guidance generation by way of ProteinGuide. $\mathbf{x}$ denotes a sequence and $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model; $y$ denotes a property of interest that we seek to guide toward using predictive model $p_{\phi}(y\mid\mathbf{x})$. $\Delta t$ denotes change in time in the diffusion process. When guiding, we blend information from the predictive model with that of the pre-trained model in a statistically correct manner. Red letters indicate mutations that are introduced when guiding with the predictive model.
@@
For example, one may want to guide an inverse-folding model to produce proteins with greater or lesser stability than the model naturally generates, or to produce proteins more likely to catalyze a particular reaction, and so forth. Perhaps more importantly, one might want to guide a model using newly acquired experimental measurements, such as catalytic activity for a desired reaction—that is, integrate experimental data in a statistically coherent manner with an existing generative model, without retraining the existing model. Such an ability would enable training-efficient iterative cycles of design, build, test, and learn—a process where experimental feedback guides subsequent AI-based design rounds.
@@
We recently showed for the first time how to correctly and tractably unlock guidance for the two specific classes of generative models of diffusion, and flow matching, on discrete state-spaces [16]. Discrete state-spaces include not only sequences of amino acid, DNA and RNA, but also SMILES strings, graphs, language tokens, structure tokens, or any object represented by a sequence of discrete tokens. Herein, we i) newly show how to use guidance on other commonly-used classes of generative models for discrete spaces, including a large class of Masked Language Models (MLMs) such as ESM3, and order-agnostic auto-regressive models (OA-AR) such as ProteinMPNN, and ii) illustrate in silico application of our method for representative use cases in protein engineering including application to ESM3 and ProteinMPNN. We call our method ProteinGuide.
@@
The essential technical formalism required to derive ProteinGuide, namely to generalize guidance to MLM and OA-AR models, is the connection between how these and diffusion models on discrete state-spaces are trained. Despite that conceptually these models seem to be quite different, it turns out that the loss functions used for training OA-AR models, a broad class of MLMs, as well as masking diffusion and flow models are equivalent (Fig. 1a and Methods). Consequently, herein, we can view a trained MLM or OA-AR model as a diffusion model on a discrete state-space. Correspondingly, we can then leverage our guidance methodology [16] for these models (Fig. 1b and Methods). Although connections between the training objectives of these model classes has been previously noted [17, 18, 19, 20, 21], the practical utility of this connection has been unclear. Here, we show that this connection enables a general framework for guiding a broad class of models. We also note that for autoregressive (AR) models, one can use FUDGE [22] for guidance, which can be viewed as a special case of ProteinGuide for the case of AR models.
@@
To illustrate the potential of ProteinGuide, we apply it to two representative, well-known protein generative models, ProteinMPNN and ESM3. For ProteinMPNN, we trained a regression model on experimental stability measurements [23] with which to guide ProteinMPNN to generate amino acid sequences that encode proteins that are more stable than what ProteinMPNN would do on its own. By employing guidance, we do so in a plug-and-play manner, rather than by retraining the inverse folding model. In our second example, we use ESM3 to generate structure tokens, so as to demonstrate the generality of ProteinGuide to not just amino acid sequences but also other sequences, such as of structure tokens. Specifically, we guide ESM3 to generate backbone structures with specified CATH fold class labels [24], by way of a classifier we trained on CATH data. Although ESM3 was trained to condition on free-form function word tokens, ProteinGuide offers a complementary, plug-and-play approach to conditioning on any property for which we have a predictive model (classifier or regressor), and so is not limited to conditioning on representations or data chosen at the time of ESM3 training.
Importantly, because our approach implements correct statistical conditioning, it preserves the knowledge embedded in the already-trained generative model, while guiding the generation process toward regions of sequence space with desired functional specification. This lies in contrast to heuristic methods such as continuing to train an existing model with new data (i.e., supervised fine-tuning), for which the statistical outcome is not well-defined, and for which important information from the already-trained model may be forgotten. Moreover, we do so without having to adjust the pre-existing model’s parameters.
Direct Preference Optimization (DPO [26, 11]) and related work enables updating of pre-trained models, although up until recently, these methods have been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models. Wallace et al. [27] extended DPO to diffusion models on continuous state-spaces by way of approximation, while Borso et al. [28] explored how to similarly do so for these models on discrete state-spaces. Importantly, the connections previously mentioned between MLMs, diffusion and flow matching models on discrete state-spaces, and order-agnostic autoregressive models, means that we can also generalize the aforementioned approximate DPO to these same model classes. Herein, we focus our study on plug-and-play guidance, rather than DPO-style approaches, although the latter could prove equally useful.
[…]
On the whole, we observed that guidance, as expected, led to desired conditioning on each of our tasks. Also as expected, when guiding ESM3 with CATH labels, more fine-grained levels and also rarer classes were more challenging—that is, had lower success rates (Fig. 2b). We expect the quality of generated structures would improve concomitantly with improved backbone generative models, independent of guidance.
@@
Figure 2: In silico demonstration of ProteinGuide with ProteinMPNN and ESM3. a. Generating structure-conditioned protein sequences from ProteinMPNN with no guidance (teal) and with ProteinGuide for stability (pink), across eight proteins (horizontal axis). The height of each bar represents the percentage of sequences predicted to be at least as stable than wildtype ($\Delta\Delta G\geq 0$), while the hatched sub-portion indicates the subset of those stable sequences that also correctly fold into the desired structure (RMSD $\leq 2$Å). b. Generating backbone structure tokens with ESM3 with no guidance (teal), and with fold class ProteinGuided ESM3 (pink), across different CATH hierarchy levels. The height of each bar shows the percentage of structures correctly reclassified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level. For level C, there are three classes: mainly $\alpha$, mainly $\beta$, and mixed $\alpha/\beta$. For level A and T, we grouped the classes for each level by their frequency of occurrence in the AFDB (Common, Regular, Rare) and calculated the metrics within each group following Geffner et al. [29]. c. Examples of randomly chosen proteins generated by guiding ESM3 with three example CATH classes.
@@ Acknowledgments
This work was funded in part by the the U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research, Lawrence Livermore National Laboratory BioSecure SFA within the Secure Biosystems Design program (SCW1710), and also the Office of Naval Research (ONR) under grant N00014-23-1-2587.
@@ Methods
[…]
@@
As an example, Campbell et al. [32] defines a forward noising process that gradually corrupts data by transitioning to a noise distribution. The model is trained to reverse this process, learning rates that can denoise corrupted samples.
There are also a number of works that extend flow matching, which were originally developed for continuous state-spaces [35, 36, 37], to discrete state-spaces [38, 39], also leveraging CTMCs, but differ from diffusion models in how the CTMCs are defined, which also induce different training objectives. These models specify a conditional flow that defines how probability mass should move from a noise distribution to the data distribution. The model learns to generate samples by following the flow trajectories. Campbell et al. [38] show that the rates used in generative sampling can be obtained by training a neural network with parameters $\theta$, $p^{\theta}(x_{1}|x_{t})$, to approximate the true denoising distribution using the standard cross-entropy loss,
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}|x_{t},t)\right],$ (2)
where $x_{1}$ is the clean data, $p(t)$, referred to as the noise schedule, is a distribution with full support on the interval $[0,1]$, and $x_{t}$ is the noised data sample from the forward noising process $p_{t|1}(x_{t}|s_{1})$. [^1] For all models, we follow the convention of flow-based models, where $t=1$ is the target (training) distribution, and $t=0$ is the noise distribution. As shown in Campbell et al. [38], the cross entropy loss can be understood as a simplification to the ELBO used to train continuous-time discrete diffusion models [32]. Given such a trained denoising model $p^{\theta}(x_{1}|x_{t})$, the unconditional rate matrix used in sampling is defined as
$R^{\theta}_{t}(x_{t},x_{t+dt})=\mathbb{E}_{x_{1}\sim p^{\theta}(x_{1}|x_{t})}\Big{[}\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})\Big{]},$ (3)
@@
A popular instantiation of both discrete state-spaces diffusion and flow models uses masking as the noising process, where tokens in a sequence are progressively replaced with mask tokens during the forward process and the model learns to recover the original tokens during the reverse process. For example, the masking conditional flow can be defined as $p^{\text{mask}}(x_{t}|x_{1})=\text{Cat}(t\delta\{x_{1},x_{t}\}+(1-t)\delta\{M,x_{t}\})$. This formulation has connections to masked language models and order-agnostic autoregressive models, which we will describe in detail next.
@@
Connections among masked discrete diffusion models, masked language models (MLM) and order-agnostic autoregressive models (OA-AR) were initially pointed out by Austin et al. [17] and Hoogeboom et al. [18]. Recent works [19, 20, 21] further clarified these connections, including the connections to CTMC-based formulation of diffusion and flow models [32, 34, 38], and showed that the continuous-time ELBO used to train masked diffusion models is essentially a weighted sum of cross entropy losses predicting the masked states from unmasked positions, where the proportion of masked positions ranges from 0 (complete masking) to 1 (no masking). For particular choices of the noise schedules, these training objectives can be shown to be equivalent to the training objective of order-agnostic autoregressive models [40], (generative) masked language models [17], or discrete flow matching models [38, 39] under a masking process. Since ESM3 [4] and ProteinMPNN [1] were trained under the MLM and OA-ARM objectives respectively, we can view them as discrete state-space diffusion and flow models, and sample from them as such, leveraging the generative rate matrices from Equation 3 induced by the trained denoising model, $p^{\theta}(x_{1}|x_{t})$. This reformulation not only provides a principled way to sample from these models, but also offers additional sampling flexibility, including, but not limited to, the usage of different noise schedules from training and corrector-sampling [38, 39], which have been shown to sometimes improve sampling results compared to prior heuristics [41] for sampling from generative masked language models [39]. In addition, this reformulation unlocks the capacity to condition the generative process via guidance, which we introduce in Section Background on diffusion guidance.

### v2

@@ Abstract
Generative machine learning models on sequences are transforming protein engineering. However, no principled framework exists for conditioning these models on auxiliary information, such as experimental data, in a plug-and-play manner. Herein, we present ProteinGuide—a principled and general method for conditioning—by unifying a broad class of protein generative models under a single framework. We demonstrate the applicability of ProteinGuide by guiding two protein generative models, ProteinMPNN and ESM3, to generate amino acid and structure token sequences, conditioned on several user-specified properties such as enhanced stability, enzyme classes, and CATH-labeled folds. We also used ProteinGuide with inverse folding models and our own experimental assay to design adenine base editor sequences for high activity.
@@
Generative models are enabling protein engineers to create bespoke proteins for therapeutic, agricultural and environmental uses [1, 2, 3]. The outcome of any AI-based protein engineering approach is to ultimately specify a protein sequence, or a set of such sequences, even when designed by way of structure. For example, one might first use a backbone structure generative model such as RFdiffusion [2], Chroma [3], or ESM3 [4], followed by use of a backbone-conditioned generative model on sequences (also called an inverse-folding model), such as ProteinMPNN [1] or ESM-IF1 [5], to generate sequences likely to fold into the specified backbones. Alternatively, one might forgo generating a backbone structure intermediary and instead generate sequences directly from an “evolutionary” sequence-only model such as ESM3 [4], ProGen [6], EvoDiff [7], DPLM [8] and others [9, 10]. In all cases, the final step employs a generative model of sequences.
@@
A key ingredient missing from the current protein engineering AI menu is the plug-and-play ability to post hoc “guide” sequence generative models toward properties of interest, because the technical underpinnings to do so have remained elusive. By guide, we mean, formally, to condition the probability distribution encoded by an existing generative model, $p(x)$, on newly-desired properties of interest, $p(y\in Y|x)$, so as to be able to generate (i.e., sample) sequences from the desired conditional distribution, $x\sim p(x|y\in Y)$. Importantly, we want to condition in a plug-and-play manner, meaning that we want to avoid re-training the existing generative model, instead modulating how we sample from it with an auxiliary model, such as a property predictive model. Although it is currently possible to re-train generative models to achieve property conditioning [11, 12], the plug-and-play characteristics of guidance would enable everyday users to readily adapt existing large, expensive-to-train models of interest to suit their needs, on-the-fly [13, 14, 15]. [^1] Herein, we use the term “guide” and “guidance” to refer to “predictor guidance” in Nisonoff et al. [16]. Interested readers can also find how to perform predictor-free guidance in that same work.
@@
Figure 1: Equivalence between classes of generative models enables guidance with ProteinGuide. a. Schematic showing how Masked Language Models (e.g., ESM3) and order-agnostic autoregressive models (e.g., ProteinMPNN) can be viewed through the lens of masking diffusion/flow models. This interpretation enables generalization of our guidance method [16] to obtain ProteinGuide. Masking of one sequence position is denoted with the symbol “?”; colored dots help to visually tag individual sequences. The sequence being masked during training is M K R S V T; $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model. The shared training objective involves randomly masking out parts of the sequence and then predicting the masked sequence given the remaining unmasked context, over different masking rates denoted by more or fewer “?” in a sequence. b. Contrasting unguided generation with guided generation by way of ProteinGuide. A sequence is denoted, $\mathbf{x}$, the pre-trained generative model is denoted $p_{\theta}(\mathbf{x})$; $y$ denotes a property of interest that we seek to guide toward using predictive model $p_{\phi}(y\mid\mathbf{x})$. When guiding, we coherently blend information from the predictive model with that of the pre-trained model in a statistically correct manner. Red letters indicate mutations that are introduced when guiding with the predictive model, and not seen when generating in an unguided manner.
@@
For example, one may want to guide an inverse-folding model to produce proteins with greater solubility than the model naturally generates, or to produce proteins more likely to bind a particular target, and so forth. Perhaps more importantly, one might want to guide a model using newly acquired experimental measurements, such as catalytic activity for a desired reaction—that is, integrate experimental data in a statistically coherent manner with an existing generative model, without re-training the existing model. Such an ability would enable training-efficient iterative cycles of design, build, test, and learn—a process where experimental feedback guides subsequent AI-based design rounds.
@@
We recently showed for the first time how to correctly and tractably perform guidance for two specific classes of generative models, diffusion and flow matching, on discrete state-spaces [16]. Discrete state-spaces include not only sequences of amino acid, DNA and RNA, but also SMILES strings, graphs, language tokens, structure tokens, or any object represented by a sequence of discrete tokens. Herein, we build on this work by i) generalizing our guidance method to two more commonly-used classes of generative models for discrete spaces, namely, a large class of Masked Language Models (MLMs) such as ESM3, and order-agnostic autoregressive models (OA-AR) such as ProteinMPNN, yielding our new method, ProteinGuide, ii) illustrating in silico application of ProteinGuide to representative use cases in protein engineering, and iii) applying ProteinGuide to obtain adenine base editors showing enhanced activity experimentally compared to the unguided model.
Figure 2: In silico and experimental demonstration of ProteinGuide. a. Generating structure-conditioned protein sequences from ProteinMPNN without guidance (teal) and with ProteinGuide for stability (pink), across eight proteins (horizontal axis). The height of each bar represents the percentage of sequences predicted to be at least as stable as wild type ($\Delta\Delta G\leq 0$), while the hatched sub-portion indicates the fraction of those stable sequences that are also predicted to correctly fold into the desired structure (RMSD $\leq 2$Å). Error bars on all panels represent standard errors of proportions, calculated as $\sqrt{p(1-p)/n}$, where $p$ is the empirical proportion of success and $n$ is the sample size. b. Using ESM3 to re-design enzyme sequences without guidance (teal) and with ProteinGuide (pink). The % residues re-designed refers to how many residues (positions) of a randomly chosen enzyme sequence from the target class were masked for re-design. When all are masked (100%), the task becomes de novo design. Bar height indicates the percentage of generated samples classified by CLEAN [17] as being from the correct (specified) target enzyme class. The hatched sub-portion shows the fraction of correctly classified samples with median pLDDT $>0.8$. c. Generating backbone structure tokens with ESM3 without guidance (teal), and with fold class ProteinGuided ESM3 (pink), across different CATH hierarchy levels. The height of each bar shows the percentage of structures correctly classified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level. For level C, there are three classes: mainly $\alpha$, mainly $\beta$, and mixed $\alpha/\beta$. For level A and T, we grouped the classes for each level by their frequency of occurrence in the AFDB (Common, Regular, Rare) and calculated the metrics within each group following Geffner et al. [18]. d. Schematic representation of the ProteinGuide ABE experiment. In round 1 we diversified a 86-amino acid TadA region using two generative models, ESM3 and FMIF, and built a library of 2,000 variants (library 1) which was assayed in a bacterial selection strain. The estimated activities of the library members were used to train a classifier to ProteinGuide for a second round of designs. This second round, library 2, contained 2,000 sequences that were cloned and assayed (round 2) and the activities of the two libraries were compared. e. A histogram of the experimentally measured log enrichments for both library 1 (teal) and library 2 (pink). ProteinGuide was used to generate the round 2 library such that variants would be enriched for having high activity (predicted activity in log enrichment units larger than the gray dashed line).
@@
The essential technical formalism required to derive ProteinGuide, namely to generalize guidance in Nisonoff et al. [16] to MLMs and OA-AR models, is the connection between how these models and diffusion models are trained. Although conceptually diffusion/flow matching models, MLMs, and OA-AR models seem to be quite different, it turns out that the loss functions used for training them are equivalent (Fig. 1a and Methods). [^2] For MLMs, the equivalence holds when the masking rate at training time covers fully masked, fully unmasked, and everything in between, such as for ESM3 (see Methods). Consequently, we can view a trained MLM or OA-AR model as a diffusion model on a discrete state-space, allowing us to leverage our earlier guidance method [16] for these models (Fig. 1b and Methods). While connections between these training objectives has been previously noted [19, 20, 21, 22, 23], the practical utility of this observation has been unclear. Herein, we show that this connection enables a general framework for guiding a broad class of models. Note that for autoregressive (AR) models, one can use FUDGE [24] for guidance, which can be viewed as a special case of ProteinGuide for the case of AR models. More detailed discussion of related work can be found in the Methods.
@@
To illustrate the potential of ProteinGuide, we applied it, in silico, to three tasks, using two representative, well-known protein generative models, ProteinMPNN and ESM3. In our first task, we trained a regression model on experimental stability measurements [25] with which to guide ProteinMPNN to generate amino acid sequences encoding proteins that are more stable than what ProteinMPNN would do on its own (Fig. 2a). By employing guidance, we do so in a plug-and-play manner, rather than by re-training the inverse folding model [11]. In our second task, we guide ESM3 to re-design enzyme sequences predicted to belong to specific enzyme classes, based on a published classifier, CLEAN, for enzyme commission number (Fig. 2b) [17]. Our re-design task spans de novo design (starting with a 100% masked sequence), to 50% re-design, wherein half the sequence is masked out at random before generation, and the rest is held constant. In our third task, we demonstrate the generality of ProteinGuide beyond amino acid sequences, to structure tokens. In particular, we guide ESM3 to generate backbone structures (as tokens) with specified CATH fold class labels [26], by way of a classifier we trained on CATH data (Fig. 2c). Although ESM3 was trained to condition on free-form function word tokens, ProteinGuide offers a complementary, plug-and-play approach to conditioning on any property for which we have a predictive model (classifier or regressor), and so is not limited to conditioning on representations or data chosen at the time of ESM3 training. Across these three tasks, we observed that guidance, as expected, led to desired conditioning (Fig. 2a-c). Also, as expected, guiding ESM3 with rare CATH classes (which includes finer-grained CATH level labels) was more challenging—that is, had lower success rates. The quality of generated structures should improve concomitantly with improved backbone generative models, independent of guidance.
Finally, we used ProteinGuide to engineer an adenine base editor for high activity (Fig. 2d). ABEs are a class of CRISPR effectors that can introduce single nucleotide changes in the DNA [27]. We started with the wild-type TadA sequence (an ABE) which is known to have very low, but detectable activity [28] and used two models, an in-house flow-matching inverse folding model, FMIF [16], and ESM3, to jointly generate one library of 2,000 variants. We next estimated the activity of these variants by their enrichment in a bacterial selection assay based on the restoration of an inactive antibiotic resistance upon base editing (see Methods). Then we trained a classifier on these data to predict high- vs. low-activity variants. Lastly, we applied ProteinGuide to design a second library of 2,000 sequences to be enriched for high activity relative to the first library. We found that the majority of the ProteinGuided variants in the second round had higher activities (mean log enrichment $0.66\pm 0.02$) than the first round library (mean log enrichment $-0.28\pm 0.01$). Furthermore, the second round library revealed variants that were more active than any seen in the first round library (Fig. 2e).
@@ Acknowledgments
This work was funded in part by the the U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research, Lawrence Livermore National Laboratory BioSecure SFA within the Secure Biosystems Design program (SCW1710), and also the Office of Naval Research (ONR) under grant N00014-23-1-2587. D.F.S. is an Investigator of the Howard Hughes Medical Institute, M.L. is a Life Science Research Fellow funded by Howard Hughes Medical Institute. For their assistance with Illumina sequencing we thank the Innovative Genomics Institute NGS facility, Netravathi Krishnappa and for nanopore and Sanger sequencing we thank the UC Berkeley DNA Sequencing Facility. We thank Hanlun Jiang, Antoine Koehl, Jianan Canal Li, and Flora Zhiqi Wang for helpful discussions.
@@ Methods
While diffusion models were initially developed for continuous state-spaces [13, 15, 29], recent work has extended these frameworks to discrete domains such as text, molecular graphs, and protein sequences. These extensions enable generative modeling of inherently discrete data without requiring continuous relaxations. There are a number of works defining discrete-time diffusion processes in discrete spaces [19, 30]. They were later extended to continuous-time [31, 32, 33] leveraging the formulation of continuous-time Markov chains (CTMCs), which describe stochastic processes where a system transitions between discrete states over continuous time. The dynamics of a CTMC are fully determined by an initial distribution $p(x_{0})$ and transition rates $R_{t}(x,\tilde{x})$ that specify the instantaneous probability of transitioning from state $x$ to state $\tilde{x}$ at time $t$:
@@
As an example, Campbell et al. [31] defines a forward noising process by specifying a time-dependent “forward” rate matrix that gradually transforms the (unknown) data distribution into a reference distribution, such as a uniform distribution over amino acid sequences. A denoising model is then trained to reverse this process, learning rates that can transform the corrupted samples back to samples from the data distribution.
There are also a number of works that extend flow matching, which were originally developed for continuous state-spaces [34, 35, 36], to discrete state-spaces [37, 38], also leveraging CTMCs, but differ from diffusion models in how the CTMCs are defined, which also induce different training objectives. These models specify a conditional flow that defines how probability mass should move from a noise distribution to the data distribution. The model learns to generate samples by following the flow trajectories. Campbell et al. [37] show that the rates used in generative sampling can be obtained by training a neural network with parameters $\theta$, $p_{\theta}(x_{1}|x_{t},t)$, to approximate the true denoising distribution using the standard cross-entropy loss,
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}|x_{t},t)\right],$ (2)
where $x_{1}$ is the clean data, $p(t)$, referred to as the noise schedule, is a distribution with full support on the interval $[0,1]$, and $x_{t}$ is the noised data sample from the forward noising process $p_{t|1}(x_{t}|x_{1})$. [^3] For all models, we follow the convention of flow-based models, where $t=1$ is the target (training) distribution, and $t=0$ is the noise distribution. As shown in Campbell et al. [37], the cross entropy loss can be understood as a simplification to the ELBO used to train continuous-time discrete diffusion models [31]. Given such a trained denoising model $p_{\theta}(x_{1}|x_{t},t)$, the (unconditional) rate matrix used in sampling is defined as
$R^{\theta}_{t}(x_{t},x_{t+dt})=\mathbb{E}_{x_{1}\sim p_{\theta}(x_{1}|x_{t})}\Big{[}\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})\Big{]},$ (3)
@@
In practice, for computational tractability, an independent forward noising process is used for each dimension, $i\in\{1,2,...,D\}$, of $x_{1}$. Campbell et al. [31, 37] show that in this setting, the reverse process also factorizes per dimension. Thus, the denoising model, $p_{\theta}(x_{1}|x_{t},t)$, need only learn a factorized distribution of the clean data and the loss becomes
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$ (4)
Similarly, the rate matrix can be expressed per dimension as
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\mathbb{E}_{x_{1}^{i}\sim p_{\theta}(x_{1}^{i}|x_{t},t)}\Big{[}\overrightarrow{R_{t}^{i}}(x_{t+dt}^{i},x_{t}^{i}|x_{1}^{i})\Big{]},$ (5)
where $\overrightarrow{R_{t}^{i}}(x_{t+dt}^{i},x_{t}^{i}|x_{1}^{i})$ is the forward noising rate matrix applied to dimension $i$.
There are many methods for sampling from discrete state-space diffusion and flow models by integrating the learned rate matrices with numerical integration schemes [31, 37, 38, 39]. One simple approach is Euler integration with a discrete time step $\Delta t$, where transition probabilities are computed as $p(x_{t+\Delta t}=\tilde{x}|x_{t}=x)=\delta_{x,\tilde{x}}+R_{t}(x,\tilde{x})\Delta t$ and new states are sampled categorically at each step. Alternative methods include Gillespie’s algorithm [40, 41] for exact simulation and $\tau$-leaping [31] for approximate sampling that allows multiple dimension to transition simultaneously. The flow matching formulation also introduces a stochasticity hyperparameter $\eta$ that can be adjusted at inference time to control the randomness of the sampling path, which Campbell et al. [37] showed could improve sample quality.
A popular instantiation of both discrete state-spaces diffusion and flow models uses masking as the per-dimension noising process, where tokens in a sequence are progressively replaced with mask tokens during the forward process and the model learns to recover the original tokens during the reverse process. For example, the masking conditional flow can be defined as $p^{\text{mask}}(x_{t}^{i}|x_{1}^{i})=\text{Cat}(t\delta\{x_{1}^{i},x_{t}^{i}\}+(1-t)\delta\{?,x_{t}^{i}\})$, where $\delta\{i,j\}$ is the Kronecker delta which is $1$ when $i=j$, 00 otherwise, and “?” denotes a mask token. Campbell et al. [37] showed that under the masking noise process, the unconditional rate matrix used in sampling defined in Equation 5 simplifies to
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\frac{p_{\theta}(x_{1}^{i}\,|\,x_{t},t)}{1-t}\delta\{x_{t}^{i},?\}.$ (6)
This formulation has connections to Masked Language Models and order-agnostic autoregressive models, which we will describe in detail next.
@@
Connections among masked discrete diffusion models, Masked Language Models (MLM) and order-agnostic autoregressive models (OA-AR) were initially pointed out by Austin et al. [19] and Hoogeboom et al. [20]. Recent works [21, 22, 23, 37, 38] further clarified these connections, including the connections to CTMC-based formulation of diffusion and flow models [31, 33, 37], and showed that the continuous-time ELBO used to train masked diffusion models is essentially a weighted sum of cross entropy losses predicting the masked states from unmasked positions, where the proportion of masked positions ranges from 0 (complete masking) to 1 (no masking). For particular choices of the noise schedules, these training objectives can be shown to be equivalent to the training objective of order-agnostic autoregressive models [42], (generative) Masked Language Models [19], or discrete flow-matching models [37, 38] under a masking process. Since ESM3 [4] and ProteinMPNN [1] were trained under the MLM and OA-AR objectives respectively, we can view them as discrete state-space diffusion and flow models, and sample from them as such, leveraging the generative rate matrices from Equation 3 induced by the trained denoising model, $p_{\theta}(x_{1}|x_{t})$. This formulation not only provides a principled way to sample from these models, but also offers additional sampling flexibility, including, but not limited to, the usage of different noise schedules during sampling from those used during training, and corrector-sampling [37, 38], which have been shown to sometimes improve sampling results compared to prior heuristics [43] for sampling from generative Masked Language Models [38]. In addition, this formulation unlocks the capacity to condition the generative process via guidance, which we introduce in a later section.